OpenStudy (anonymous):

find the next number in the pattern 1,8,27,64,125,....

OpenStudy (mertsj):

It look like those are all cubes of integers.

OpenStudy (btaylor):

the rule is \[a_n = n^3\] so the 6th term would be 6^3 = 216.

OpenStudy (ajprincess):

1^3=1 2^3=8 3^3=27 4^3=64 5^3=125 6^3=216